When rolling a dice what is your chance of getting three?
1 in 6. As there are 6 different numbers. 1/6 as a fraction
How many 4 digit numbers can you make that contain 5?
There is a clever but tricky way to solve this. If we have to find the number of 4-digit integers that contains at least one 5, we can also find the number of all the 4-digit integers and the number of integers that do not contain any 5's and subtract it from the first number. This is called complementary counting.
So, first of all, we must find the number of 4-digit numbers there are. There are 9000 of them.
Now, we find the number of 4-digit integers without 5's. By thinking a little bit, we see the first digit must be from 0~9 excluding 5. That is a total of 9 numbers. This is the same for the next three digits. Therefore, there are 94 = 6561 4-digit numbers without a 5.
Finally, we can subtract. 9000 - 6561 = 2439
Why is the normal probability distribution widely used in practice?
Suppose you have a random variable, X, with any distribution. Suppose you take a sample of n independent observations, X1, X2, ... Xn and calculate their mean. Repeat this process several times. Then as the sample size increases and the number of repeats increases, the distribution of the means tends towards a normal distribution. This is due to the Central Limit Theorem.
One consequence is that many common statistical measures have an approximately normal distribution.
What is the probability of rolling a number that is divisible by 3 or a 5?
If you roll the die often enough the probability is near enough 1 so as to make no difference.
On a single roll, the probability is 1/2.
How many different ways can 6 people stand in line at a cash register?
First place can be any of 6, second any of 5 etc. 6 x 5 x 4 x 3 x 2 = 720 different ways!
The probability of rolling an even die and then an odd die is (3 in 6) times (3 in 6), which is the same as (1 in 2) times (1 in 2), which is the same as 0.5 time 0.5, which is 0.25.
Twins are not necessarily identical. They can be fraternal as well. Identical twins result when one fertilized egg splits in two. Fraternal twins result from two fertilized eggs.
What is the ratio of 4 blue marbles to 6 white marbles?
The ratio of 4 blue marbles to 6 white marbles can be said as 4:6, 4 to 6, or 4/6.
This can then be simplified to 2:3 by a factor of 2.
What is the probability of rolling an odd number on a 20 sided dice?
Assuming the sides are labelled one to twenty and that all sides are the same shsp and size and that the dice in not biased in any way, then there will be a 50% chance of getting an odd number.
(4/52)*(3/51)*(2/50) = 24/132600 = 0.000181
What is a run of five cards of different suits?
A run of 5 cards, all sequential ( Ace, 2, 3, 4, 5, different suits) for example, in poker is known as a straight.
If these sequential runs of cards are of the same suit, then they are known as a straight flush.
What is the probability of 1 6 4 on 3 rolls of a die?
What is the probability of 1, 6, 4, on 3 rolls of a die
It is 13.2%
Why is an increase in the number of varieties of a good regarded as a gain from trade?
More varieties provide buyers with more options. More options allow buyers to better match their needs and costs, thus achieve more value for their money.
"Dice" is both singular and plural. Usually, it's one "die" and two "dice."
Two six sided dice are rolled which number has the lowest probability of being rolled?
If two six sided fair dice are rolled, the sum of the result of both dice that has the
lowest probability to come up is 2 and 12. P(2) = 1/36. P(12) = 1/36.
When tossing 5 coins simultaneously what is the probability that at least 1 head is showing?
Probability of no heads = (0.5)^5 = 0.03125
Probability of at least one head = 1 - probability of no heads = 1 - 0.03125 = 0.96875
How many ways can 5 friend sit together at the movies in 5 seats?
120 ways if ABCDE is considered different from EDCBA. In these two cases, the same people are next to one another , only the left and right re changed.
The probability of getting a red chip on the first draw is 5/15.
The probability of getting a red chip on the second draw, given that the first draw
gave a red chip that was not replaced in the bag is 4/14.
The probability that the first and second draw give a red chip is ;
5/15 x 4/14 = 2/21 ≈ 0.0952 = 9.52 %